# Beyond Formula Complexity: Effective Information Criterion Improves Performance and Interpretability for Symbolic Regression

### v1

@@ 5 Conclusion and Discussion
We propose the Effective Information Criterion (EIC), a measure of structural plausibility for symbolic regression formulas with interpretations from numerical analysis and information-processing perspectives. Unlike complexity-based metrics, EIC detects unreasonable structures that affect interpretability. It improves heuristic search-based SR by steering away from unreasonable formulas, enhances generative methods by filtering implausible samples, and aligns well with expert judgments for interpretability. Despite its effectiveness, several avenues remain for future work. First, integrating EIC guidance with pre-trained model guidance may further boost performance. Second, analyzing EIC at each subformula, rather than only the maximum, could more precisely identify redundant or unreasonable structures, enabling finer-grained evaluation. Finally, in complex systems where simple microscopic rules produce emergent macroscopic behaviors, such as phase transitions, it remains to be studied whether the resulting dynamics formulas retain low EIC scores.
Appendix A Methodology Details
A.1 Effective Information Criterion
Using equation 2 and equation 3, we can derive that:
EIC $=N-M$ (7)
[…]
$=\log_{10}(\delta_{r}^{2}/\sigma_{r}^{2}),$
demonstrating the physical meaning of EIC from the perspective of the information-processing systems. Based on this equation, we give the calculation algorithm of EIC as in Algorithm 1.
Input: Formula $f$, input data $X$, hyperparameter $\sigma_{r}^{2}$ ($\sigma_{r}^{2}\rightarrow 0$)
Output: Noisy output $\tilde{y}$, clean output $y$, EIC value
1
2if $f$ is an unary operator then
 $\text{operator}\leftarrow f.\text{operator}$ ;
 // $\text{operator}\in\{\sin,\cos,\tan,\exp,\log,\cdots\}$
3 $\text{operand}\leftarrow f.\text{operand}$;
4 $(\tilde{x},x,\text{EIC})\leftarrow\text{CalculateEIC}(\text{operand},X)$;
5 $y\leftarrow\text{operator}(x)$;
[…]
7 $\tilde{y}\leftarrow\tilde{y}+\epsilon\tilde{y},\ \epsilon\sim\mathcal{N}(0,\sigma_{r}^{2})$;
8 $\delta_{r}^{2}\leftarrow\mathrm{Var}\left[\frac{\tilde{y}-y}{y}\right]$;
@@
9 $\text{EIC}\leftarrow\max\left\{\text{EIC},\log_{10}(\delta_{r}^{2}/\sigma_{r}^{2})\right\}$;
10 return $(\tilde{y},y,\text{EIC})$;
11
12 else if $f$ is a binary operator then
 $\text{operator}\leftarrow f.\text{operator}$ ;
 // $\text{operator}\in\{+,-,\times,\div,\cdots\}$
13 $\text{operand}_{1}\leftarrow f.\text{operand}[1]$;
14 $\text{operand}_{2}\leftarrow f.\text{operand}[2]$;
15 $(\tilde{x}_{1},x_{1},\text{EIC}_{1})\leftarrow\text{CalculateEIC}(\text{operand}_{1},X)$;
16 $(\tilde{x}_{2},x_{2},\text{EIC}_{2})\leftarrow\text{CalculateEIC}(\text{operand}_{2},X)$;
17 $y\leftarrow f(x_{1},x_{2})$;
18 $\tilde{y}\leftarrow f(\tilde{x}_{1},\tilde{x}_{2})$;
19 $\tilde{y}\leftarrow\tilde{y}+\epsilon\tilde{y},\ \epsilon\sim\mathcal{N}(0,\sigma_{r}^{2})$;
20 $\delta_{r}^{2}\leftarrow\mathrm{Var}\left[\frac{\tilde{y}-y}{y}\right]$;
21 $\text{EIC}\leftarrow\max\left\{\text{EIC}_{1},\text{EIC}_{2},\log_{10}(\delta_{r}^{2}/\sigma_{r}^{2})\right\}$;
22 return $(\tilde{y},y,\text{EIC})$;
23
24 else
 return $(f(X),f(X),0)$ ;
 // $f$ is a variable or a constant
25
Algorithm 1 CalculateEIC
In this research, we kept the hyperparameter $\sigma_{r}$ as $10^{-6}$. We also find that, when the $\sigma_{r}$ is small enough, the value of EIC is independent of the specific value of $\sigma_{r}^{2}$ or $N$. This actually makes our EIC a parameter-free metric, reflecting the intrinsic plausibility of a formula independent of external tuning.
A.2 Enhance Search-based Methods with EIC

### v2

@@ 5 Conclusion and Discussion
We propose the Effective Information Criterion (EIC), a metric to quantify the structural stability of symbolic regression formulas. EIC measures the amplification of inherent rounding noise during recursive calculation, offering a grounded interpretation from both numerical precision and signal-processing perspectives. Unlike traditional complexity-based metrics, EIC effectively evaluates structural rationality and detects ill-conditioned structures that render formulas physically implausible. Our results demonstrate that integrating EIC can steer search-based SR methods away from irrational formulas to yield superior Pareto frontiers and can boost the sample efficiency of generative SR methods by filtering implausible samples. Furthermore, EIC aligns with human expert preference for interpretability in 70% of cases, validating structural stability as a critical prerequisite for human-perceived interpretability. Despite its effectiveness, several avenues remain for future work. First, integrating EIC guidance with pre-trained model guidance may further boost performance. Second, analyzing EIC at each subformula, rather than only the maximum, may more precisely identify redundant or unreasonable structures, enabling finer-grained evaluation. Finally, in complex systems where simple microscopic rules produce emergent macroscopic behaviors, such as phase transitions, it remains to be studied whether the resulting dynamics formulas retain low EIC scores.
Impact Statement
This paper presents the Effective Information Criterion (EIC), a metric designed to quantify the structural rationality of formulas discovered by symbolic regression algorithms. By prioritizing formulas with structural stability, our work contributes to more reliable AI-driven scientific discovery. We do not foresee any direct negative social impacts from this research.
Appendix A Effective Information Criterion (EIC)
A.1 Proof of Proposition 3.2
Proof.
First, we show that the cumulated error $\eta_{k}$ is a zero-mean random variable with a variance that tends to 0 as $\sigma\to 0$.
Lemma A.1.
Under a first-order approximation, $\eta_{k}$ is a zero-mean random variable given by the recursive linearization:
$\eta_{k}\approx\epsilon_{k}+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}\times\eta_{i},$ (8)
where $\kappa_{k,i}\triangleq\frac{y_{i}}{y_{k}}\frac{\partial e_{k}}{\partial y_{i}}$ is the partial relative condition number of the operation at node $k$ with respect to input $i$. The approximation becomes exact in the limit $\sigma\to 0$.
Proof.
Let the inputs to node $k$ be $y_{1},\dots,y_{i},\dots,y_{m}$, which correspond to the outputs of child nodes. For the leaf nodes, we have $\tilde{y}_{k}=y_{k}(1+\epsilon_{k})$ and $\eta_{k}=(\tilde{y}_{k}-y_{k})/y_{k}$, so $\eta_{k}=\epsilon_{k}$ for leafe nodes, whose mean is zero and whose variance, $\sigma$, tends to zero when $\sigma\to 0$. Then, according to Assumption 3.1, we have
$\tilde{y}_{k}=e_{k}(\tilde{y}_{1},\dots,\tilde{y}_{m})\times(1+\epsilon_{k})$ (9)
Substituting $\tilde{y}_{i}$:
$\tilde{y}_{k}=e_{k}(y_{1}\times(1+\eta_{1}),\dots,y_{m}\times(1+\eta_{m}))\times(1+\epsilon_{k})$ (10)
We perform a first-order Taylor expansion of $e$ around the noise-free values $\mathbf{y}=(y_{1},\dots,y_{m})$, obtaining
$e_{k}(y_{1}\times(1+\eta_{1}),\dots,y_{m}\times(1+\eta_{m}))\approx e_{k}(\mathbf{y})+\sum_{i=1}^{m}\frac{\partial e_{k}}{\partial y_{i}}(y_{i}\eta_{i})$ (11)
Substituting this back into the expression for $\tilde{y}_{k}$, and noting that $y_{k}=e_{k}(\mathbf{y})$:
$\tilde{y}_{k}\approx\left(y_{k}+\sum_{i=1}^{m}\frac{\partial e_{k}}{\partial y_{i}}y_{i}\eta_{i}\right)\times(1+\epsilon_{k})$ (12)
Expanding the product and ignoring the second-order error term ($\eta_{i}\cdot\epsilon_{k}\ll\min\{\eta_{i},\epsilon_{k}\}$, since $|\eta_{i}|\ll 1~\text{and}~|\epsilon_{k}|\ll 1~\text{when}~\sigma\rightarrow 0$):
$\tilde{y}_{k}\approx y_{k}+y_{k}\epsilon_{k}+\sum_{i=1}^{m}\frac{\partial e_{k}}{\partial y_{i}}y_{i}\eta_{i}$ (13)
Dividing both sides by $y_{k}$, we identify the accumulated relative error $\eta_{k}\triangleq(\tilde{y}_{k}-y_{k})/y_{k}$ as:
$\eta_{k}\approx\epsilon_{k}+\sum_{i=1}^{m}\underbrace{\left(\frac{y_{i}}{y_{k}}\frac{\partial e_{k}}{\partial y_{i}}\right)}_{\kappa_{k,i}}\eta_{i}$ (14)
Since $\epsilon_{k}$ has zero mean and is independent of $\eta_{i}$ (which are linear combinations of zero-mean noise from previous steps and is thus independent of each other), the expectation $\mathbb{E}[\eta_{k}]$ is a linear combination of zeros, satisfying $\mathbb{E}[\eta_{k}]=0$, and its variance $\mathrm{Var}[\eta_{k}]\rightarrow 0$ when $\sigma\rightarrow 0$ (since both $\epsilon_{k}$ and $\eta_{i}$ $\rightarrow 0$), which further suggests $|\eta_{k}|\rightarrow 0$ when $\sigma\rightarrow 0$. This confirms that the cumulated error $\eta_{k}$ is a zero-mean random variable whose variance tends to zero when $\sigma\to 0$. ∎
Then, we can prove the Eq (3) by applying the conditional variance operator $\text{Var}[\cdot|x]$ to both sides in Eq (8). Two key independence properties apply here:
  1. 1.
    $\epsilon_{k}$ is the intrinsic noise introduced at the current node and is independent of the propagated errors $\{\eta_{i}\}_{i\in\mathcal{C}[k]}$ from previous steps.
  2. 2.
    Since the formula is structured as a computational tree, the sub-expressions corresponding to different children are disjoint. Thus, the accumulated errors $\eta_{i}$ from different children are mutually independent random variables.
Using these independence properties, the variance of the sum becomes the sum of the variances:
$\text{Var}[\eta_{k}|x]\approx\text{Var}[\epsilon_{k}]+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}^{2}(x)\text{Var}[\eta_{i}|x],$ (15)
where the approximation becomes exact in the limit $\sigma\to 0$.
Substituting $\text{Var}[\epsilon_{k}]=\sigma^{2}$, dividing both sides by $\sigma^{2}$, and taking the limit $\sigma\to 0$ (which makes the linear approximation exact), we get
$\lim_{\sigma\to 0}\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}}=1+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}^{2}(x)\left(\lim_{\sigma\to 0}\frac{\text{Var}[\eta_{i}|x]}{\sigma^{2}}\right)$ (16)
which provides the recursive relation to conclude the proof:
$s_{k}^{2}(x)=1+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}^{2}(x)s_{i}^{2}(x)$ (17)
A.2 Remark for the definition of EIC
In Eq (4), we define EIC as the maximum value of the logarithm $\bar{s}_{k}$ of all nodes $k\in\mathcal{T}[f]$, rather than the logarithm of the root node $\bar{s}_{\text{root}}$. The choice of the maximum operator is motivated by the non-monotonic nature of error propagation. Observing the recurrence relation in eq (3), if the squared relative condition number is small ($\kappa_{k,i}^{2}\ll 1$), the cumulative relative error $s_{k}(x)$ does not necessarily increase relative to its children’s $S_{i}$. Physically, a vanishing $\kappa_{k,i}$ implies that the child node $i$ exerts negligible influence on its parent $k$. In symbolic regression, this indicates structural redundancy, that is, the existence of components that do not contribute to describing the data, which violates the principle of Occam’s Razor [8]. Crucially, such a redundant sub-structure $i$ can be numerically ill-conditioned (exhibiting a very high $s_{i}(x)$), yet its instability is masked by the small $\kappa_{k,i}$, resulting in a deceptively low $s_{\text{root}}(x)$. By defining EIC as the maximum over all nodes ($\max_{k}$), we explicitly detect and penalize these redundant yet ill-conditioned internal structures, ensuring that a physically plausible formula implies stability not just at the output, but throughout its entire computational graph.
A.3 Proof of Eq (5)
Proof.
Based on the definition of EIC and $\bar{s}_{k},s_{k}(x)$, we have
EIC $\triangleq\max_{k\in\mathcal{T}[f]}\log_{10}\bar{s}_{k}$ (18)
$\triangleq\max_{k\in\mathcal{T}[f]}\log_{10}\sqrt{\mathbb{E}_{x\sim\mathcal{D}}[s_{k}^{2}(x)]}$
$=\frac{1}{2}\max_{k\in\mathcal{T}[f]}\log_{10}\mathbb{E}_{x\sim\mathcal{D}}[s_{k}^{2}(x)]$
@@ A.3 Proof of Eq (5)
$\triangleq\frac{1}{2}\max_{k\in\mathcal{T}[f]}\log_{10}\mathbb{E}_{x\sim\mathcal{D}}\left[\lim_{\sigma\to 0}\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}}\right]$
Here, the limit operator can be sequentially commuted with the expectation, logarithm, and maximum operators due to the linearity of expectation, the continuity of the logarithm, and the finiteness of the node set $\mathcal{T}[f]$, respectively:
EIC $=\frac{1}{2}\max_{k\in\mathcal{T}[f]}\log_{10}\lim_{\sigma\to 0}\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}}\right]$ (19)
$=\frac{1}{2}\max_{k\in\mathcal{T}[f]}\lim_{\sigma\to 0}\log_{10}\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}}\right]$
$=\frac{1}{2}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}}\right]$
Finally, to merge the conditional variance and expectation into the global variance, consider law of total variance decomposition:
$\mathbb{E}_{x}[\text{Var}[\eta_{k}|x]]=\text{Var}[\eta_{k}]-\text{Var}_{x}[\mathbb{E}[\eta_{k}|x]],$ (20)
where the second term is a higher-order infinitesimal of the first term when $\sigma\to 0$: consider that the conditional mean relative error is $\mathbb{E}[\eta_{k}|x]\approx 0$. The non-zero bias arises only from second-order terms (e.g., $\eta^{2}$), implying $\mathbb{E}[\eta_{k}|x]=O(\sigma^{2})$. Consequently, the variance of this bias scales as the square of the order:
$\text{Var}_{x}[\mathbb{E}[\eta_{k}|x]]=\text{Var}[O(\sigma^{2})]=O(\sigma^{4}).$ (21)
Substituting this back, the fraction inside the limit becomes:
$\frac{\mathbb{E}_{x}[\text{Var}[\eta_{k}|x]]}{\sigma^{2}}=\frac{\text{Var}[\eta_{k}]-O(\sigma^{4})}{\sigma^{2}}=\frac{\text{Var}[\eta_{k}]}{\sigma^{2}}-\underbrace{O(\sigma^{2})}_{\to 0}.$ (22)
Thus, in the limit $\sigma\to 0$, the bias term vanishes, allowing us to replace the expected conditional variance with the global variance:
$\mathbb{E}_{x}[\text{Var}[\eta_{k}|x]]=\text{Var}[\eta_{k}]\quad(\sigma\to 0).$ (23)
Substitude this into Eq (19), we finally prove that
EIC $=\frac{1}{2}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\text{Var}[\eta_{k}]}{\sigma^{2}}$ (24)
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\sqrt{\frac{\text{Var}[\eta_{k}]}{\sigma^{2}}}$
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\sqrt{\text{Var}[\eta_{k}]}}{\sqrt{\text{Var}[\epsilon_{k}]}}$
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{{\text{Std}[\eta_{k}]}}{{\text{Std}[\epsilon_{k}]}}.$
A.4 Invariance of EIC Under Unit Transformation
Proof.
Here we prove that the EIC is invariant under physical unit transformations (e.g., rescaling input variables from meters to centimeters). This invariance arises from the fact that the partial relative condition number $\kappa$ is intrinsically dimensionless, rendering the subsequent $s_{k}(x)$ and EIC metric independent of the chosen system of units.
Specifically, let $[x]$ denote the physical dimension of a variable $x$ (e.g., Length, Time). For any computational operation $y_{k}=e(\{y_{i}\}_{i\in\mathcal{C}[k]})$, the partial derivative $\frac{\partial y_{k}}{\partial y_{i}}$ represents the rate of change of the output with respect to the input, and thus carries the dimension $y_{k}][y_{i}]^{-1}$. The normalization term $\frac{y_{i}}{y_{k}}$ carries the dimension $y_{i}][y_{k}]^{-1}$.
Multiplying these terms to obtain $\kappa_{k,i}$:
$\kappa_{k,i}]=\left[\frac{y_{i}}{y_{k}}\right]\cdot\left[\frac{\partial y_{k}}{\partial y_{i}}\right]=([y_{i}][y_{k}]^{-1})\cdot([y_{k}][y_{i}]^{-1})=1.$ (25)
This confirms that $\kappa_{k,i}$ is a dimensionless quantity (a pure number).
Physical unit transformations (e.g., $x\to\lambda x$) act as linear scalings on the numerical values of dimensional quantities. However, dimensionless quantities are, by definition, invariant to such scaling because the scaling factors in the numerator and denominator cancel out exactly. Thus, the value of $\kappa_{k,i}$ remains constant regardless of the unit system.
As established in Proposition 3.2, the pointwise variance amplification $s_{k}(x)$ is computed recursively using solely the invariant $\kappa$ terms and the constant $1$. Consequently, both $s_{k}(x)$ and the final EIC derived from it are invariant under physical unit transformations. ∎
A.5 Calculation of EIC
Here, we use a simple example to illustrate how EIC is calculated and demonstrate its ability to assess the structural stability of formulas. Specifically, consider the identity $f(x)=x$ versus its mathematically equivalent but redundant form $g(x)=\ln(\exp(x))$. For the concise form $f(x)$, the pointwise amplification is constantly $s_{f}(x)=1$. Thus, we have $\bar{s}_{f}=1$ and $\text{EIC}(f)=\log_{10}1=0$, suggesting no cumulated information loss during the calculation. For the redundant form $g(x)$, however, the intermediate node $u=\exp(x)$ has a condition number $\kappa_{u,x}=x$, leading to $s_{u}^{2}(x)=1+x^{2}$. The final node $y=\ln(u)$ has $\kappa_{y,u}=1/x$, yielding $s_{y}^{2}(x)=1+(1/x)^{2}(1+x^{2})=2+x^{-2}$. When inputs $x$ follow a distribution with a large mean (e.g., $x\sim\mathcal{N}(\mu,1)$ with $|\mu|\gg 1$), the intermediate node $u$ dominates the instability ($\bar{s}_{u}^{2}=\mathbb{E}[s_{u}^{2}(x)]=2+\mu^{2}\gg\bar{s}_{y}^{2}=\mathbb{E}[s_{y}^{2}(x)]=2+\mathbb{E}[x^{-2}]$), yielding:
$\text{EIC}(g)=\log_{10}\bar{s}_{u}=\log_{10}\sqrt{2+\mu^{2}}\approx\log_{10}\mu,$ (26)
which is larger than $\text{EIC}(f)=0$ since $\mu\gg 1$. This result aligns with intuition: when inputs are distributed in a numerically large region, the exponentiation operation induces severe structural instability. Even though the subsequent logarithm restores the value to the correct range, the precision of the intermediate result remains hypersensitive to microscopic perturbations in $x$. This example highlights that EIC assesses a formula not merely as a functional mapping (Input $\to$ Output), but as an information processing system. By evaluating the stability of the entire computational graph, EIC discerns the structural rationality of the representation, allowing it to identify the superior, robust realization, even among mathematically equivalent candidates.
This computation process can be formally described by pseudocode Algorithm 1. It is worth noting that although Algorithm 1 requires a differential engine capable of calculating conditional partial derivatives, EIC can be estimated based on Eq (5). Specifically, this estimation strategy chooses a small enough $\sigma$ and traverses the symbolic tree in post-order, maintaining two parallel computation paths to simulate the noise propagation process: a “clean” reference path computed using standard double-precision arithmetic, and a “noisy” perturbed path where multiplicative noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ is injected at every operation node. Consequently, the instability factor $\bar{s}_{k}$ is estimated via the variance of the relative error between these two paths, which is then mapped to the EIC value according to Eq. (4). While this stochastic approach shares the same time complexity $O(N)$ as the analytical method (where $N$ is the number of nodes), it is significantly easier to integrate into existing evaluation pipelines as it requires only standard forward passes. We detailedly describe this estimation strategy in Algorithm 2.
Input: Formula $f$ and data distribution $\mathcal{D}$ with NN samples.
Output: Variance amplification factor $s(x)$ and output value $y(x)$ at root node, as well as EIC value of $f$.
$k\leftarrow f.\text{root}$  % Start from the root node
if $k$ is a Leaf Node (Variable or Constant) then
  $y_{k}(x)\leftarrow\text{Evaluate}(k,\mathcal{D})$  % Get values of $k$ in $\mathcal{D}$
  $s_{k}^{2}(x)\leftarrow 1$
  $\bar{s}_{k}^{2}\leftarrow\mathbb{E}_{x\sim\mathcal{D}}[s_{k}^{2}(x)]$  % $\bar{s}_{k}=1$ for leaf nodes
  $\text{EIC}_{k}\leftarrow\log_{10}\bar{s}_{k}$  % $\text{EIC}=0$ for leaf nodes
  return $(s_{k}(x),y_{k}(x),\text{EIC}_{k})$
else
  $\text{EIC}_{k}\leftarrow 0$
  $e_{k}\leftarrow k.\text{operator}$
  for each child $i\in\mathcal{C}[k]$ do
   $f^{\prime}\leftarrow\text{Subtree rooted at }i$
   $(s_{i}(x),y_{i}(x),\text{EIC}_{i})\leftarrow\text{CalculateEIC}(f^{\prime},\mathcal{D})$
   $\text{EIC}_{k}\leftarrow\max\{\text{EIC}_{k},\text{EIC}_{i}\}$
  end for
  $y_{k}(x)\leftarrow e_{k}(\{y_{i}\}_{i\in\mathcal{C}[k]})$
  for each child $i\in\mathcal{C}[k]$ do
   $\kappa_{k,i}(x)\leftarrow\frac{y_{i}(x)}{y_{k}(x)}\frac{\partial e_{k}}{\partial y_{i}}|_{x\sim\mathcal{D}}$
  end for
  $s_{k}^{2}(x)\leftarrow 1+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}^{2}(x)s_{i}^{2}(x)$
  $\bar{s}_{k}^{2}\leftarrow\mathbb{E}_{x\sim\mathbb{E}}[s_{k}^{2}(x)]$
  $\text{EIC}_{k}\leftarrow\max\{\text{EIC}_{k},\log_{10}\bar{s}_{k}\}$
  return $(s_{k}(x),y_{k}(x),\text{EIC}_{k})$
end if
Algorithm 1 CalculateEIC (Recursively calculate partial relative condition number)
Input: Formula $f$ and data distribution $\mathcal{D}$ with NN samples.
Parameter: A small enough noise scale $\sigma$.
Output: Noisy output $\tilde{y}(x)$ and clean output $y(x)$ at root node, as well as EIC value of $f$.
$k\leftarrow f.\text{root}$  % Start from the root node
if $k$ is a Leaf Node (Variable or Constant) then
  $y_{k}(x)\leftarrow\text{Evaluate}(k,\mathcal{D})$  % Get values of $k$ in $\mathcal{D}$
  Generate $\epsilon_{k}\sim\mathcal{N}(0,\sigma^{2})$ of size NN
  $\tilde{y}_{k}(x)\leftarrow y_{k}(x)\times(1+\epsilon_{k})$
  $\eta_{k}\leftarrow\frac{\tilde{y}_{k}(x)-y_{k}(x)}{y_{k}(x)}$  % $\eta_{k}=\epsilon_{k}$ for leaf nodes
  $\text{EIC}_{k}\leftarrow\log_{10}\sqrt{\text{Var}[\eta_{k}]/\sigma^{2}}$  % $\text{EIC}=0$ for leaf nodes
  return $(\tilde{y}_{k}(x),y_{k}(x),\text{EIC}_{k})$
else
  $\text{EIC}_{k}\leftarrow 0$
  $e_{k}\leftarrow k.\text{operator}$
  for each child $i\in\mathcal{C}[k]$ do
   $f^{\prime}\leftarrow\text{Subtree rooted at }i$
   $(\tilde{y}_{i}(x),y_{i}(x),\text{EIC}_{i})\leftarrow\text{EstimateEIC}(f^{\prime},\mathcal{D},\sigma)$
   $\text{EIC}_{k}\leftarrow\max\{\text{EIC}_{k},\text{EIC}_{i}\}$
  end for
  $y_{k}(x)\leftarrow e_{k}(\{y_{i}\}_{i\in\mathcal{C}[k]})$
  Generate $\epsilon_{k}\sim\mathcal{N}(0,\sigma^{2})$ of size NN
  $\tilde{y}_{k}(x)\leftarrow e_{k}(\{\tilde{y}_{i}\}_{i\in\mathcal{C}[k]})\times(1+\epsilon_{k})$
  $\eta_{k}\leftarrow\frac{\tilde{y}_{k}(x)-y_{k}(x)}{y_{k}(x)}$
  $\text{EIC}_{k}\leftarrow\max\{\text{EIC}_{k},\log_{10}\sqrt{\text{Var}[\eta_{k}]/\sigma^{2}}\}$ % Eq. (5)
  return $(\tilde{y}_{k}(x),y_{k}(x),\text{EIC}_{k})$
end if
Algorithm 2 EstimateEIC (Recursively add stochastic perturbation)
A.6 Interpretations for EIC
Numerical Precision Perspective. EIC measures the stability of a formula’s computational structure when multiplicative noise of a given intensity is introduced, which, from the perspective of numerical precision, corresponds to a specified number of significant digits.
According to quantization theory, the operation of rounding a value to retain $N$ significant decimal digits is statistically equivalent to injecting a multiplicative noise with variance $\sigma^{2}\approx 10^{-2N}$, or $\sigma\approx 10^{-N}$ [30]. Based on this equivalence, the injected noise $\epsilon_{k}$ with standard error $\text{Std}[\epsilon_{k}]=\sigma$ in Assumption 3.1 can be considered as the operation of rounding with a precision of $N\approx-\log_{10}(\text{Std}[\epsilon_{k}])$ significant digits.  [^3] Although the rounding error typically follows a Uniform distribution, our derivation relies solely on the second moment. Therefore, the number of digits $N$ (and $M_{k}$) discussed here is defined in the variance-equivalent sense, mapping the noise intensity to the precision level that would produce such variance regardless of the specific underlying distribution. On the other hand, the accumulated relative (i.e., multiplicative) error $\eta_{k}$ implies that only $M_{k}\approx-\log_{10}(\text{Std}[\eta_{k}])$ effective significant digits remain at node $k$ due to the accumulation of injected noise. Using this relationship, we can formally link EIC to the loss of precision. Specifically, in the high-precision limit ($\sigma\to 0$, corresponding to calculation precision $N\to\infty$), while the effective precision $M_{k}$ at any node $k$ increases with $N$ indefinitely, it asymptotically lags behind $N$ by a constant margin. We define the maximum inherent precision loss induced by the computational structure as:
$\Delta_{\max}\triangleq\lim_{N\to\infty}(N-\min_{k\in\mathcal{T}}M_{k}).$ (27)
which represents the maximum asymptotic gap between the applied computational precision and the effective precision achievable at any node $K$. Based on the relationship between variance and precision, the EIC equals this loss:
$\text{EIC}=\Delta_{\max}.$ (28)
This indicates that the EIC measures the structural instability in terms of lost significant digits.
Proof.
To proof this, we can substitute the relationship $N\approx-\log_{10}(\text{Std}[\epsilon_{k}])$ and $M_{k}\approx-\log_{10}(\text{Std}[\eta_{k}])$ into Eq (5), obtaining:
EIC $=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\text{Std}[\eta_{k}]}{\text{Std}[\epsilon_{k}]}$ (29)
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}(\log_{10}\text{Std}[\eta_{k}]-\log_{10}\text{Std}[\epsilon_{k}])$
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}}((-M_{k})-(-N))$
$=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}}(N-M_{k})$
$=\lim_{\sigma\to 0}(N-\min_{k\in\mathcal{T}}M_{k})$
This completes the proof. ∎
Signal Processing Perspective. EIC can alternatively be interpreted through the perspective of signal processing. By treating the computational graph as a signal transmission channel, we can analyze how the quality of the “information signal” degrades as it propagates from input to output.
For any node $k$, we define the “signal” as the noise-free output value $y_{k}$, and the “noise” as the deviation caused by the accumulated relative error $\tilde{y}_{k}-y_{k}=y_{k}\eta_{k}$. The local Signal-to-Noise Ratio (SNR) is defined as the ratio of signal power to noise power:
$\text{SNR}_{k}\triangleq\frac{\text{Signal Power}}{\text{Noise Power}}=\frac{y_{k}^{2}}{\text{Var}[y_{k}\eta_{k}]}=\frac{y_{k}^{2}}{y_{k}^{2}\text{Var}[\eta_{k}]}=\frac{1}{\text{Var}[\eta_{k}]}.$ (30)
Similarly, the intrinsic SNR is.
$\text{SNR}_{\text{intrinsic}}\triangleq\frac{y_{k}^{2}}{\text{Var}[y_{k}\epsilon_{k}]}=\frac{1}{\sigma^{2}}$ (31)
To quantify the degradation of signal quality, we utilize the Noise Figure $F_{k}\triangleq 10\log_{10}\frac{\text{SNR}_{\text{intrinsic}}}{\text{SNR}_{k}}$, which measures the ratio of input SNR to output SNR.
From this perspective, we can interpret EIC as the noise figure (scaled by $\frac{1}{20}$ at the bottleneck nodes:
$\text{EIC}=\frac{1}{20}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}F_{k}$ (32)
Proof.
Begin from the Eq (5), we have
EIC $=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\text{Std}[\eta_{k}]}{\text{Std}[\epsilon_{k}]}$ (33)
$=\frac{1}{2}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\text{Var}[\eta_{k}]}{\sigma^{2}}$
$=\frac{1}{2}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\text{SNR}_{\text{intrinsic}}}{\text{SNR}_{k}}$
$=\frac{1}{20}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}F_{k}.$
This completes the proof. ∎
Appendix B Methodology Details
B.1 Enhance Search-based Methods with EIC